OpenStudy (lgbasallote):

\[y = \sin (\cot x)\] okay my question here is just...do you simplify first to \(\cos x\) then derivative? or derivative first then simplify?

OpenStudy (ganpat):

i guess, straightway derive ...

OpenStudy (lgbasallote):

so it's product rule right?

OpenStudy (lgbasallote):

ohh lol i thought it was sinx cotx

OpenStudy (lgbasallote):

cot x is the argument..

OpenStudy (ganpat):

derive is at sinx and then , derive cot x as well..

OpenStudy (lgbasallote):

so it's \[\cos (\cot x) \times \csc^2 x?\]

OpenStudy (lgbasallote):

\[\csc^2 x \cos (\cot x)\]

OpenStudy (anonymous):

\[y=\sin(f(x))\] \[y'=\sin'((f(x))*f'(x)\]

OpenStudy (lgbasallote):

so...im right right?

OpenStudy (anonymous):

\[\cos(cotx)*\cot'(x)\]

OpenStudy (lgbasallote):

...so im right? lol

OpenStudy (ganpat):

\[\cos (\cot x) * - cosec ^{2} x \]

OpenStudy (lgbasallote):

negative?

OpenStudy (ganpat):

right..

OpenStudy (anonymous):

\[-\cos(cotx)*\csc^2x\]

OpenStudy (lgbasallote):

oh wait yeah...derivative of cofunction is negative

OpenStudy (lgbasallote):

thanks guys

OpenStudy (ganpat):

w/c..